OpenStudy (anonymous):

Simplify the given equation. 17x - 6 + 3x - 5 = x + 11 + 4x 20x - 11 = 5x + 11 20x + 1 = 5x + 11 9x = 16x

OpenStudy (anonymous):

Doing it step by step... 17x - 6 + 3x - 5 = x + 11 + 4x 20x-11=5x+11 15x=22 x=22/15

OpenStudy (anonymous):

The answer would be 20x-6+3x-5= x+11+4x

OpenStudy (anonymous):

I mean 20x-11= 5x+11 lol

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome